if an isosceles triangle has two angles measuring 45 degrees and 67.5 degress what is the measure of the third triangle
Answer by rwm(914)

isosceles triangle means two angels are the same size.
so it is either 45 or 67.5
it can't be 45 because 45 +45=90
67.5 is not 90
67.5 +67.5=135
135+45=180
we could also have gone x=180-45-67.5